Lasky Company sold merchandise with a list price of USD 60,000 on July 1. For each of the

following independent assumptions, calculate (1) the gross selling price used to record the sale and (2) the amount

that the buyer would have to remit when paying the invoice.

Trade Discount Credit Date

Granted Terms Paid

a. 30%, 20% 2/10, n/30 July 10

b. 40%, 10% 2/EOM, n/60 August 10

c. 30%, 10%, 5% 3/10/EOM, n/60 August 10

d. 40% 1/10, n/30 July 12


1
Expert's answer
2021-07-12T16:18:21-0400

Solution:

a.

(i) Determination of gross selling price (Price before cash discount/net price)

First trade discount = 30% = $(30% "\times" 60,000) = $18000

First net price = List price - First trade discount = $(60000 - 18000) = $42000

Second trade discount = 20% "\times" First net price = $(20%"\times"42000) = $8400

Second (final) net price = First net price - Second trade discount = $(42000 - 8400) = $33600


(ii) Determination of price remitted by the customer (cash price):

cash price = Final net price - cash discount

Cash discount = n/30, therefore no cash discount was accorded, thus the cash price is equal to net price = $33600


b.

(i) Determination of gross selling price (Price before cash discount/net price)

First trade discount = $(40% "\times" 60,000) = $24000

First net price = List price - First trade discount = $(60000 - 24000) = $36000

Second trade discount = 10% "\times" First net price = $(10%"\times"36000) = $3600

Second (final) net price = First net price - Second trade discount = $(36000 - 3600) = $32400


(ii) Determination of price remitted by the customer (cash price):

Cash price = Final net price - Cash discount

Cash discount = 2/EOM, payment was made on August 10(still within 20 days after the 26th), therefore cash discount = 2% of the final net price = $(2%"\times"32400) = $3240

cash price = Net price - Cash discount = $(32400-3240) = $29160


c.

(i)Determination of gross selling price (Price before cash discount/net price)

First trade discount = $(30% "\times" 60,000) = 18000

First net price = List price - First trade discount = $(60000 - 18000) = $42000

Second trade discount = 10% "\times" First net price = $(10%"\times"42000) = $4200

Second net price = First net price - Second trade discount = $(42000 - 4200) = $37800

Third trade discount = 5% of second net price = $(5%"\times"37800) = $1890

Final net price = second net price - third trade discount = $(37800-1890) = $35910


(ii) Determination of price remitted by the customer (cash price):

Cash price = Final net price - cash discount

Cash discount = 2/10/EOM, payment was made on August 10 (still within 20 days after the 26th), therefore cash discount = 3% of the final net price = $(3%"\times"35910) = $1077.30

Cash price = Net price - Cash discount = $(35910-1077.30) = $34832.70


d.

(i)Determination of gross selling price (Price before cash discount/Net price)

First trade discount = $(40% "\times" 60,000) = $24000

Net price = List price - Trade discount = $(60000 - 24000) = $36000


(ii) Determination of price remitted by the customer (cash price):

Cash price = Final net price - cash discount

Cash discount = 1% if payment is made within 10 days = 1% of net price = $(0.01"\times"36000) = $360

"\therefore" Cash price = $(36000-360) = 35640